# Reduce Computational Complexity For Continuous Wavelet Transform In Acoustic Recognition Using Hop Size

### v1

@@ Abstract
Index Terms:
Continuous Wavelet Transform, Hop Size, Acoustic Recognition.
@@ I Introduction
Wavelet Transform (WT) is increasingly applied to acoustic recognition tasks due to its multiresolution analysis capability, which enhances the performance of trained models [1]. Numerous researchers have contributed significantly to this field. Copiaco et al.[2] employed the scalogram of the Continuous Wavelet Transform (CWT) as a spectro-temporal feature extractor for domestic audio classification. The scalogram images were fed into a model comprising Convolutional Neural Networks (CNNs) and a Support Vector Machine (SVM) for the classification task. Their research on the DCASE 2018 Task 5 dataset demonstrated a substantial improvement compared to top-performing models. Gupta, Chodingala, and Patil[3] utilized CNNs as the prediction model and scalograms generated from CWT as input features for voice liveness detection. Their method effectively distinguished between live speech and spoofing attempts. They proposed a handcrafted Morlet wavelet, achieving a prediction performance of 80% accuracy, surpassing the conventional Short-Time Fourier Transform (STFT) spectrogram, which achieved 62.08% accuracy. Chatterjee et al.[4] developed an approach for musical instrument identification employing a combination of Convolutional Siamese Network and Residual Siamese Network as the deep learning model. Audio excerpts were transformed into scalograms as time-frequency features using CWT. The method achieved a high classification accuracy of 80% for different instrumental audio from public datasets with only five training datasets. Phan[5] conducted a performance comparison between CWT and STFT as inputs for a CNNs prediction model. The results showed the advantage of CWT over STFT in recognizing non-stationary machine noise. This research also highlighted a drawback of CWT due to its high computational complexity. As evidenced by the recent publications mentioned above, CWT is widely used as a temporal-spectral feature extractor. However, the computational expense for CWT is significant, as it is computed continuously for every sample of a discrete signal, generating a large amount of data that may contain redundancy due to the similarity of adjacent data samples. Therefore, an approach that exploits the benefits of multiresolution analysis of CWT to enhance the prediction performance of trained models while maintaining low computational expense is highly anticipated.
@@ II-A Wavelet transform
WT is a technique that decomposes a signal into a form that better represents the original signal’s features for further processing [6]. In acoustic recognition, WT converts a one-dimensional (1D) time signal into a two-dimensional (2D) time-frequency plane, as described by the calculation formula in (1). WT is a function of time translation b and frequency shift a. The signal’s energy is normalized by the factor 1/$\sqrt{\textit{a}}$ to ensure consistent energy levels across all frequency scales. The wavelet is contracted and dilated according to the varying scale, and each scaled wavelet is then shifted along the time axis to convolve with the signal x(t).
@@
On the other hand, the Discrete WT (DWT) is computed by discretizing the values of scale and translation according to powers of 2, which is why it is often referred to as the dyadic wavelet transform [1]. Unlike the CWT, the DWT is not directly implemented through the inner product of the original signal and the wavelet function. Instead, it is realized using a filter bank followed by down-sampling. The signal is decomposed up to level m with scale $\textit{a}=2^{m}$ and translation $\textit{b}=n2^{m}$. This type of transform is computationally more efficient than the CWT. However, the result of the transformation is not a matrix of coefficients, as the number of coefficients is reduced by half at each scale, making it unsuitable for generating heat maps in acoustic recognition task. The scalogram of the DWT is illustrated in Fig. 1(c), where the generated coefficients are reconstructed to form a matrix. In comparison with the CWT, the DWT scalogram exhibits lower time-frequency resolution, often appearing as a fragmented image. Conversely, the CWT scalogram is more condensed and contains richer information regarding the time and frequency characteristics of the signal.
@@ III-A Dataset
The experiment utilizes the MIMII dataset [7], which contains real-world sound data from factory environments, including sounds from fans, pumps, sliders, and valves. For each machine type, there are two categories of sound: normal sounds, representing machines operating correctly, and abnormal sounds, indicating faulty machinery. The objective of this dataset is to train a model capable of detecting faulty machines. The recorded audio is mixed with background noise at three different signal-to-noise ratio (SNR) levels: -6 dB, 0 dB, and 6 dB. The dataset comprises a total of 54,507 audio files, each 10 seconds in length, sampled at a rate of 16 kHz, resulting in 160,000 samples per file.
@@ III-B Experimental design
The research employs the PyWavelets library [8] for wavelet transformation, the TensorFlow library [9] for implementing the binary classification CNNs model, and the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [10] as the prediction performance metric. The hop size $H$ is set to 128, as in preliminary training sessions, this is considered a suitable compromise between the prediction performance of the trained model and the reduction of computational load. A study [11], which employed Mel-frequency cepstral coefficients (MFCC) as a feature extractor, and conducted the same classification task on the dataset, is used as a benchmark to assess the effectiveness of the developed methods.
@@ III-C Results
The prediction performance of the models across various audio types is documented in Tables I, II, III and IV, and visualized in Figures 3, 4, 5 and 6. As observed, the prediction capability of all models improves with increasing SNR levels. While the baseline model demonstrates comparable or superior performance to WT in the cases of fan, pump, and slider audio, the opposite trend is observed in valve audio, where WT with multi-resolution analysis outperforms MFCC with linear resolution, consistent with the findings of a previous study [5]. The performance of models indicates that the developed methods achieved satisfactory results.

### v2

@@ I Introduction
Wavelet Transform (WT) is increasingly applied to acoustic recognition tasks due to its multiresolution analysis capability, which enhances the performance of trained models [guo2022review]. Numerous researchers have contributed significantly to this field. Copiaco et al.[copiaco2019scalogram] employed the scalogram of the Continuous Wavelet Transform (CWT) as a spectro-temporal feature extractor for domestic audio classification. The scalogram images were fed into a model comprising Convolutional Neural Networks (CNNs) and a Support Vector Machine (SVM) for the classification task. Their research on the DCASE 2018 Task 5 dataset demonstrated a substantial improvement compared to top-performing models. Gupta, Chodingala, and Patil[gupta2022morlet] utilized CNNs as the prediction model and scalograms generated from CWT as input features for voice liveness detection. Their method effectively distinguished between live speech and spoofing attempts. They proposed a handcrafted Morlet wavelet, achieving a prediction performance of 80% accuracy, surpassing the conventional Short-Time Fourier Transform (STFT) spectrogram, which achieved 62.08% accuracy. Chatterjee et al.[chatterjee2023deep] developed an approach for musical instrument identification employing a combination of Convolutional Siamese Network and Residual Siamese Network as the deep learning model. Audio excerpts were transformed into scalograms as time-frequency features using CWT. The method achieved a high classification accuracy of 80% for different instrumental audio from public datasets with only five training datasets. Phan[phan2024comparison] conducted a performance comparison between CWT and STFT as inputs for a CNNs prediction model. The results showed the advantage of CWT over STFT in recognizing non-stationary machine noise. This research also highlighted a drawback of CWT due to its high computational complexity. As evidenced by the recent publications mentioned above, CWT is widely used as a temporal-spectral feature extractor. However, the computational expense for CWT is significant, as it is computed continuously for every sample of a discrete signal, generating a large amount of data that may contain redundancy due to the similarity of adjacent data samples. Therefore, an approach that exploits the benefits of multiresolution analysis of CWT to enhance the prediction performance of trained models while maintaining low computational expense is highly anticipated.
@@ II-A Wavelet transform
$X_{WT}(\textit{b},\textit{a})=\frac{1}{\sqrt{\textit{a}}}\int_{-\infty}^{\infty}\,x(t)\psi^{*}(\frac{t-\textit{b}}{\textit{a}})\,dt$ (1)
@@
On the other hand, the Discrete WT (DWT) is computed by discretizing the values of scale and translation according to powers of 2, which is why it is often referred to as the dyadic wavelet transform [guo2022review]. Unlike the CWT, the DWT is not directly implemented through the inner product of the original signal and the wavelet function. Instead, it is realized using a filter bank followed by down-sampling. The signal is decomposed up to level m with scale $\textit{a}=2^{m}$ and translation $\textit{b}=n2^{m}$. This type of transform is computationally more efficient than the CWT. However, the result of the transformation is not a matrix of coefficients, as the number of coefficients is reduced by half at each scale, making it unsuitable for generating heat maps in acoustic recognition task. The scalogram of the DWT is illustrated in Fig. 1(c), where the generated coefficients are reconstructed to form a matrix. In comparison with the CWT, the DWT scalogram exhibits lower time-frequency resolution, often appearing as a fragmented image. Conversely, the CWT scalogram is more condensed and contains richer information regarding the time and frequency characteristics of the signal.
@@ III-A Dataset
The experiment utilizes the MIMII dataset [purohit2019mimii], which contains real-world sound data from factory environments, including sounds from fans, pumps, sliders, and valves. For each machine type, there are two categories of sound: normal sounds, representing machines operating correctly, and abnormal sounds, indicating faulty machinery. The objective of this dataset is to train a model capable of detecting faulty machines. The recorded audio is mixed with background noise at three different signal-to-noise ratio (SNR) levels: -6 dB, 0 dB, and 6 dB. The dataset comprises a total of 54,507 audio files, each 10 seconds in length, sampled at a rate of 16 kHz, resulting in 160,000 samples per file.
@@ III-B Experimental design
The research employs the PyWavelets library [lee2019pywavelets] for wavelet transformation, the TensorFlow library [tensorflow2015-whitepaper] for implementing the binary classification CNNs model, and the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [provost1998case] as the prediction performance metric. The hop size $H$ is set to 128, as in preliminary training sessions, this is considered a suitable compromise between the prediction performance of the trained model and the reduction of computational load. A study [gantert2021supervised], which employed Mel-frequency cepstral coefficients (MFCC) as a feature extractor, and conducted the same classification task on the dataset, is used as a benchmark to assess the effectiveness of the developed methods.
@@ III-C Results
The prediction performance of the models across various audio types is documented in Tables I, II, III and IV, and visualized in Figures 3, 4, 5 and 6. As observed, the prediction capability of all models improves with increasing SNR levels. While the baseline model demonstrates comparable or superior performance to WT in the cases of fan, pump, and slider audio, the opposite trend is observed in valve audio, where WT with multi-resolution analysis outperforms MFCC with linear resolution, consistent with the findings of a previous study [phan2024comparison]. The performance of models indicates that the developed methods achieved satisfactory results.